# Position Paper: An Integrated Perspective on Data, Metrics, and Methodology for Deep Time-Series Forecasting

### v3

@@ 3 The ProbTS Tool
4 Analyses
Model Exchange Rate Solar Electricity Traffic Wikipedia
CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE
Glob. mean $0.188$ $0.188$ $1.403$ $1.403$ $0.412$ $0.412$ $0.540$ $0.540$ $0.577$ $0.577$
Batch mean $0.012$ $0.012$ $1.244$ $1.244$ $0.365$ $0.365$ $0.503$ $0.503$ $0.336$ $0.336$
Table 3: Results ($\textrm{mean}_{\textrm{std}}$) on long-term forecasting scenarios, each containing five independent runs with different seeds. The input sequence length is set to 36 for the ILI dataset and 96 for the others. Due to the excessive time and memory consumption of CSDI in producing long-term forecasts, its results are unavailable in some datasets.
Model pred DLinear PatchTST GRU NVP TimeGrad CSDI
len CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE CRPS NMAE
ETTm1 96 $0.282_{.002}$ $0.282_{.002}$ $0.272_{.001}$ $0.272_{.001}$ $0.383_{.053}$ $0.488_{.058}$ $0.522_{.105}$ $0.645_{.129}$ $0.236_{.006}$ $0.308_{.005}$
192 $0.309_{.004}$ $0.309_{.004}$ $0.295_{.001}$ $0.295_{.001}$ $0.396_{.030}$ $0.514_{.042}$ $0.603_{.092}$ $0.748_{.084}$ $0.291_{.025}$ $0.377_{.026}$
336 $0.338_{.008}$ $0.338_{.008}$ $0.323_{.001}$ $0.323_{.001}$ $0.486_{.032}$ $0.630_{.029}$ $0.601_{.028}$ $0.759_{.015}$ $0.322_{.033}$ $0.419_{.042}$
720 $0.387_{.006}$ $0.387_{.006}$ $0.353_{.001}$ $0.353_{.001}$ $0.546_{.036}$ $0.707_{.050}$ $0.621_{.037}$ $0.793_{.034}$ $0.448_{.038}$ $0.578_{.051}$
ETTm2 96 $0.138_{.000}$ $0.138_{.000}$ $0.132_{.001}$ $0.132_{.001}$ $0.319_{.044}$ $0.413_{.059}$ $0.427_{.042}$ $0.525_{.047}$ $0.115_{.009}$ $0.146_{.012}$
192 $0.163_{.003}$ $0.163_{.003}$ $0.157_{.001}$ $0.157_{.001}$ $0.326_{.025}$ $0.427_{.033}$ $0.424_{.061}$ $0.530_{.060}$ $0.147_{.008}$ $0.189_{.012}$
336 $0.188_{.001}$ $0.188_{.001}$ $0.176_{.000}$ $0.176_{.000}$ $0.449_{.145}$ $0.580_{.169}$ $0.469_{.049}$ $0.566_{.047}$ $0.190_{.018}$ $0.248_{.024}$
720 $0.219_{.003}$ $0.219_{.003}$ $0.205_{.001}$ $0.205_{.001}$ $0.561_{.273}$ $0.749_{.385}$ $0.470_{.054}$ $0.561_{.044}$ $0.239_{.035}$ $0.306_{.040}$
ETTh1 96 $0.352_{.011}$ $0.352_{.011}$ $0.328_{.003}$ $0.328_{.003}$ $0.379_{.030}$ $0.481_{.037}$ $0.455_{.046}$ $0.585_{.058}$ $0.437_{.018}$ $0.557_{.022}$
192 $0.393_{.001}$ $0.393_{.001}$ $0.359_{.002}$ $0.359_{.002}$ $0.425_{.019}$ $0.531_{.018}$ $0.516_{.038}$ $0.680_{.058}$ $0.496_{.051}$ $0.625_{.065}$
336 $0.419_{.007}$ $0.419_{.007}$ $0.384_{.002}$ $0.384_{.002}$ $0.458_{.054}$ $0.580_{.064}$ $0.512_{.026}$ $0.666_{.047}$ $0.454_{.025}$ $0.574_{.026}$
720 $0.502_{.029}$ $0.502_{.029}$ $0.397_{.002}$ $0.397_{.002}$ $0.502_{.039}$ $0.643_{.046}$ $0.523_{.027}$ $0.672_{.015}$ $0.528_{.012}$ $0.657_{.014}$
ETTh2 96 $0.211_{.027}$ $0.211_{.027}$ $0.177_{.000}$ $0.177_{.000}$ $0.432_{.141}$ $0.548_{.158}$ $0.358_{.026}$ $0.448_{.031}$ $0.164_{.013}$ $0.214_{.018}$
192 $0.238_{.028}$ $0.238_{.028}$ $0.201_{.001}$ $0.201_{.001}$ $0.625_{.170}$ $0.766_{.223}$ $0.457_{.081}$ $0.575_{.089}$ $0.226_{.018}$ $0.294_{.027}$
336 $0.284_{.008}$ $0.284_{.008}$ $0.240_{.001}$ $0.240_{.001}$ $0.793_{.319}$ $0.942_{.408}$ $0.481_{.078}$ $0.606_{.095}$ $0.274_{.022}$ $0.353_{.028}$
720 $0.307_{.000}$ $0.307_{.000}$ $0.252_{.000}$ $0.252_{.000}$ $0.539_{.090}$ $0.688_{.161}$ $0.445_{.016}$ $0.550_{.018}$ $0.302_{.040}$ $0.382_{.030}$
Electricity 96 $0.090_{.001}$ $0.090_{.001}$ $0.086_{.001}$ $0.086_{.001}$ $0.094_{.003}$ $0.118_{.003}$ $0.096_{.002}$ $0.119_{.003}$ $0.153_{.137}$ $0.203_{.189}$
192 $0.095_{.001}$ $0.095_{.001}$ $0.092_{.001}$ $0.092_{.001}$ $0.097_{.002}$ $0.121_{.003}$ $0.100_{.004}$ $0.124_{.005}$ $0.200_{.094}$ $0.264_{.129}$
336 $0.104_{.000}$ $0.104_{.000}$ $0.100_{.000}$ $0.100_{.000}$ $0.099_{.001}$ $0.123_{.001}$ $0.102_{.007}$ $0.126_{.008}$
720 $0.122_{.001}$ $0.122_{.001}$ $0.116_{.000}$ $0.116_{.000}$ $0.114_{.013}$ $0.144_{.017}$ $0.108_{.003}$ $0.134_{.004}$
Traffic 96 $0.356_{.009}$ $0.356_{.009}$ $0.248_{.001}$ $0.248_{.001}$ $0.187_{.002}$ $0.231_{.003}$ $0.202_{.004}$ $0.234_{.006}$
192 $0.346_{.009}$ $0.346_{.009}$ $0.245_{.001}$ $0.245_{.001}$ $0.192_{.001}$ $0.236_{.002}$ $0.208_{.003}$ $0.239_{.004}$
336 $0.350_{.008}$ $0.350_{.008}$ $0.257_{.002}$ $0.257_{.002}$ $0.201_{.004}$ $0.248_{.006}$ $0.213_{.003}$ $0.246_{.003}$
720 $0.365_{.009}$ $0.365_{.009}$ $0.266_{.001}$ $0.266_{.001}$ $0.211_{.004}$ $0.264_{.006}$ $0.220_{.002}$ $0.263_{.001}$
Weather 96 $0.112_{.001}$ $0.112_{.001}$ $0.087_{.002}$ $0.087_{.002}$ $0.116_{.013}$ $0.145_{.017}$ $0.130_{.017}$ $0.164_{.023}$ $0.068_{.008}$ $0.087_{.012}$
192 $0.122_{.001}$ $0.122_{.001}$ $0.090_{.001}$ $0.090_{.001}$ $0.122_{.021}$ $0.147_{.025}$ $0.127_{.019}$ $0.158_{.024}$ $0.068_{.006}$ $0.086_{.007}$
336 $0.130_{.002}$ $0.130_{.002}$ $0.092_{.002}$ $0.092_{.002}$ $0.128_{.011}$ $0.160_{.012}$ $0.130_{.006}$ $0.162_{.006}$ $0.083_{.002}$ $0.098_{.002}$
720 $0.144_{.001}$ $0.144_{.001}$ $0.094_{.001}$ $0.094_{.001}$ $0.110_{.004}$ $0.135_{.008}$ $0.113_{.011}$ $0.136_{.020}$ $0.087_{.003}$ $0.102_{.005}$
Exchange 96 $0.024_{.000}$ $0.024_{.000}$ $0.023_{.000}$ $0.023_{.000}$ $0.071_{.006}$ $0.091_{.009}$ $0.068_{.003}$ $0.079_{.002}$ $0.028_{.003}$ $0.036_{.005}$
192 $0.035_{.000}$ $0.035_{.000}$ $0.034_{.000}$ $0.034_{.000}$ $0.068_{.004}$ $0.087_{.005}$ $0.087_{.013}$ $0.100_{.019}$ $0.045_{.003}$ $0.058_{.005}$
336 $0.048_{.001}$ $0.048_{.001}$ $0.048_{.000}$ $0.048_{.000}$ $0.072_{.002}$ $0.091_{.002}$ $0.074_{.009}$ $0.086_{.008}$ $0.060_{.004}$ $0.076_{.006}$
720 $0.075_{.002}$ $0.075_{.002}$ $0.072_{.000}$ $0.072_{.000}$ $0.079_{.009}$ $0.103_{.009}$ $0.099_{.015}$ $0.113_{.016}$ $0.143_{.020}$ $0.173_{.020}$
ILI 24 $0.213_{.038}$ $0.213_{.038}$ $0.169_{.005}$ $0.169_{.005}$ $0.257_{.003}$ $0.283_{.001}$ $0.275_{.047}$ $0.296_{.044}$ $0.250_{.013}$ $0.263_{.012}$
36 $0.230_{.015}$ $0.230_{.015}$ $0.156_{.005}$ $0.156_{.005}$ $0.281_{.004}$ $0.307_{.007}$ $0.272_{.057}$ $0.298_{.048}$ $0.285_{.010}$ $0.298_{.011}$
48 $0.221_{.009}$ $0.221_{.009}$ $0.156_{.008}$ $0.156_{.008}$ $0.288_{.008}$ $0.314_{.009}$ $0.295_{.033}$ $0.320_{.025}$ $0.285_{.036}$ $0.301_{.034}$
60 $0.230_{.013}$ $0.230_{.013}$ $0.147_{.003}$ $0.147_{.003}$ $0.307_{.005}$ $0.333_{.005}$ $0.295_{.083}$ $0.325_{.068}$ $0.283_{.012}$ $0.299_{.013}$
[Long-term forecasting (NMAE).] Refer to caption [Long-term forecasting (CRPS).] Refer to caption     Refer to caption
[Short-term forecasting (NMAE).] Refer to caption [Short-term forecasting (CRPS).] Refer to caption     Refer to caption
Figure 2: We present plots of the metrics against non-Gaussianity for various models across all forecasting scenarios. In each subplot, a vertical line of points represents one forecasting scenario, denoted by a sequence number at the top. We distinguish between point and probabilistic forecasting methods using different colors, and we use unique markers to indicate various models.
@@ 4 Analyses
[CRPS w.r.t. seasonality strength.] Refer to caption     [The CRPS Gap, computed as TimeGrad.CRPS - CSDI.CRPS, w.r.t. trend and seasonal strengths.] Refer to caption
Figure 3: We analyze CRPS in relation to various factors to scrutinize the strengths and weaknesses of autoregressive probabilistic methods. Subplot (a), based on ETTm and ILI datasets, illustrates the error amplification that occurs with increased prediction length. Subplots (b) and (c) encompass all long-term forecasting scenarios, aiming to explore the relationship between CRPS and the strengths of trends and seasonality, respectively. Subplot (d) includes all forecasting scenarios, with a lighter and smaller circle denoting a preference for autoregressive methods over non-autoregressive ones.
Utilizing ProbTS, we benchmarked selected methods in both short-term and long-term forecasting scenarios. We evaluated these methods using both point and distributional forecast metrics. To ensure the reliability of our results and minimize randomness, we performed each experiment five times with different seeds for parameter initialization. The results for short-term and long-term forecasting are presented in Table 2 and Table 3, respectively.
Selected Methods for Comparison.
We have chosen representative methods from various research branches for comparison. Our selection is based on a comprehensive consideration of factors such as state-of-the-art performance, reproducibility, as well as the simplicity, universality, and novelty of the method designs. Adhering to these principles, we have selected PatchTST (Nie et al., 2023), TimesNet (Wu et al., 2023), N-HiTS (Challu et al., 2023), LTSF-Linear (NLinear and DLinear) (Zeng et al., 2023) from the research line that concentrates on developing customized neural architectures for long-term point forecasting. In addition, we have chosen advanced probabilistic forecasting methods for comparison, including those built upon normalizing flows such as GRU NVP, GRU MAF, and Trans MAF (Rasul et al., 2021b), and diffusion models such as TimeGrad (Rasul et al., 2021a) and CSDI (Tashiro et al., 2021). We also include general network architectures, such as Linear, GRU (Chung et al., 2014), and Transformer (Vaswani et al., 2017), to demonstrate the additional benefits brought by advanced architecture designs or probabilistic estimations. Simple non-parametric baselines, such as calculating the global mean or batch mean as all forecasts, are also included to assess the predictability of a forecasting scenario.
4.1 Discrepancies in Point and Distributional Metrics.
Our analysis revealed that optimal performance in distributional metrics does not guarantee optimal performance in point forecasts. This pattern emerged in several instances, notably in Table 3. For example, CSDI excels in the CRPS metric but falls short in NMAE across multiple datasets, such as ETTm1, ETTm2, and ETTh2. These observations underscore the necessity of a comprehensive evaluation utilizing both NMAE and CRPS metrics to gain a holistic understanding of performance in both point and distributional forecasting. It is worth highlighting that these two metrics become equivalent for point forecasting methods when a point forecast is converted into a degenerate distribution, centered on the forecasted point, for CRPS evaluation (Gneiting & Raftery, 2007).
4.2 Examination of Customized Neural Architectures
Our findings affirm the progress in this area of research. For instance, as observed in Table 3, the state-of-the-art architecture, PatchTST, consistently surpasses other models, especially when assessed using the NMAE metric. The only exception is on the Traffic dataset, where PatchTST is outperformed by GRU NVP and TimeGrad. This discrepancy is primarily attributable to the different decoding schemes utilized, which will be elaborated in Section 4.3.
@@ 4.2 Examination of Customized Neural Architectures
Diminishing Advantages of Customized Architectures in Short-term Forecasting Scenarios.
Interestingly, the significant superiority of PatchTST diminishes in short-term forecasting scenarios. This trend is evident when comparing Table 3 and 2. More intuitively, Figures 22 display the relative strengths of point and probabilistic forecasting methods in terms of NMAE, and correlates their NMAE values with our quantified indicator, non-Gaussianity, used for assessing data distribution complexity. We further observe that as non-Gaussianity increases, signifying a more complex data distribution, a noticeable performance gap emerges. This suggests that the reducing superiority of these customized architectures may be due to their limited capability to handle complex data distributions.
Challenges in Integrating Advanced Distribution Estimation Capabilities with These Architectures.
Regrettably, the integration of these architectures with advanced distribution estimation capabilities remains a challenging task. As will be discussed subsequently, non-autoregressive probabilistic methods encounter significant efficiency issues in long-term forecasting. Intriguingly, all these architectures select a non-autoregressive decoding scheme. We will delve into the advantages and disadvantages of autoregressive and non-autoregressive schemes in the next section.
4.3 Assessment of Probabilistic Forecasting Methods
At the outset, we validate the efficacy of these probabilistic methods in the short-term forecasting scenarios they were designed for. Evidence from Table 2 demonstrates that innovative methods such as CSDI and TimeGrad yield superior results in most cases.
However, in Long-term Forecasting Scenarios, Probabilistic Methods Fall Short, Even When Evaluated Using Distributional Metrics.
As depicted in Table 3, in a multitude of long-term forecasting scenarios, customized neural architectures, created specifically with point forecasting objectives in mind, decidedly outperform probabilistic methods, even when evaluated using distributional metrics such as CRPS. This suggests that existing probabilistic methods may encounter significant challenges when applied to long-term forecasting, leaving substantial room for improvement in this domain. To further illustrate this, we compare the CRPS performance of these models with the non-Gaussianity in both long-term and short-term scenarios, as depicted in Figure 2 and 2, respectively. While non-Gaussianity does provide some insights—its values are relatively lower in long-term scenarios compared to short-term contexts—these explanations are not comprehensive. This is because it remains unclear to what extent a customized neural model, coupled with effective probabilistic estimations, can perform. To delve deeper, we further examine other unique aspects of the existing probabilistic forecasting methods in search of potential clues.
The Dichotomy of Decoding Scheme Preferences.
Upon further investigation, we find that these probabilistic forecasting methods exhibit a balanced preference for both autoregressive and non-autoregressive decoding schemes. For instance, TimeGrad employs an autoregressive decoding scheme, whereas CSDI utilizes a non-autoregressive decoding approach. This contrasts starkly with the aforementioned customized architectures, which solely opt for non-autoregressive decoding. These two types of decoding schemes, however, confront distinctive challenges when applied to long-term probabilistic forecasting.
  • Autoregressive probabilistic methods, for example, TimeGrad, are prone to propagate more errors as the forecasting horizon increases or the trend strength enlarges. This phenomenon is evident in Figure 3 and 3, where we observe that the performance gap escalates with either extended prediction lengths or intensified trend strengths.
@@ 4.3 Assessment of Probabilistic Forecasting Methods
  • In contrast, non-autoregressive probabilistic methods, like CSDI, struggle with significant efficiency challenges, both in memory usage and time consumption, in long-term forecasting scenarios. This issue is further detailed in Appendix Section B.6. Furthermore, as indicated in Table 3, even for relatively small-scale datasets, such as ETTm2 and ETTh1, CSDI’s performance in long-term scenarios is less than optimal. This suggests that its learning efficacy is also compromised by the extension of the forecasting horizon.
The Unexpected Superiority of Autoregressive Decoding Scheme in Addressing Strong Seasonality.
Interestingly, the autoregressive decoding scheme exhibits a remarkable proficiency in dealing with strong seasonality. A prime example is the results for the Traffic dataset shown in Table 3, where both GRU NVP and TimeGrad outperform PatchTST, despite utilizing relatively standard neural architectures without any customization. In Figure 3, we further investigate this phenomenon by relating the strength of seasonality to the variance in performance gaps. Additionally, in Figure 3, we conduct a collective analysis of the interactions among CRPS Gap, trend, and seasonality strengths, where the CRPS Gap is defined as the difference between the CRPS scores of TimeGrad (autoregressive) and CSDI (non-autoregressive). A clear pattern emerges: as the strength of seasonality heightens, autoregressive methods gain more advantages. This observation suggests that despite its challenges, the autoregressive decoding scheme should not be dismissed in long-term forecasting. Its unique affinity with seasonality, possibly due to superior parameter efficiency, may regain significance if we can develop effective strategies to mitigate the issue of error propagation.
4.4 Future Directions
Drawing on our analyses, we propose several promising directions for future research. Firstly, the development of hybrid learning objectives that optimize both point and distributional forecasts represents a significant advancement. This strategy could foster models capable of delivering more comprehensive predictions, catering to a diverse range of practical applications that necessitate both accurate predictions and well-quantified uncertainties. Secondly, there is an urgent need for the effective integration of neural architecture design with advanced probabilistic estimation paradigms. Addressing this demand could lay the groundwork for more efficient and precise forecasting models. Particularly, overcoming the efficiency challenges associated with non-autoregressive decoding schemes and reducing error propagation in autoregressive probabilistic models, specifically in the presence of marked trends, could substantially enhance the models’ performance and practicality. Lastly, the exploration of the necessity and potential designs of customized architectures tailored for short-term forecasting could initiate a novel research field. Such exploration necessitates consideration of data characteristics beyond trends and seasonality.
4.5 Limitations
Despite the thoroughness of our study, there are several limitations. First, due to the vast number of methods available in this field, we could not include all of them in our analysis. Our selection is based on a set of criteria, but inevitably, some methods were left out. Second, while we have tried to incorporate a diverse range of datasets, there are certainly more datasets beyond those included in our study. To mitigate this limitation, we have conducted additional experiments on univariate time-series datasets, which are included in Appendix B.3. Finally, our study is primarily empirical in nature. While we have made significant efforts to conduct rigorous statistical analyses, we have not provided theoretical proofs for our results.

### v4

@@ 3 The ProbTS Tool
4 Results and Analyses
[Long-term Forecasting (NMAE).] Refer to caption [Long-term Forecasting (CRPS).] Refer to caption
[Short-term Forecasting (NMAE).] Refer to caption [Short-term Forecasting (CRPS).] Refer to caption
Figure 2: We present a comprehensive comparison between classical models designed for long-term point forecasting and short-term distributional forecasting across various prediction horizons. It utilizes a non-Gaussianity score to highlight the complexity of the data distribution across different datasets. The aggregated performance metrics are derived from Tables 9 and 10.
@@ 4 Results and Analyses
[…]
Figure 3: We explore the challenges faced by current models in conducting long-term distributional forecasting, with insights drawn from Table 10. Subplot (a) highlights the substantial error increases in AR-based probabilistic models (e.g., GRU NVP, TimeGrad) when applied to ETTm and ILI datasets. Subplots (b) and (c) examine how the strengths of trends and seasonality, respectively, affect the overall performance across all long-term forecasting datasets and horizons. Subplot (d) further investigates the combined effects of trend and seasonality, using lighter and smaller circles to indicate situations where AR-based models are favored over NAR-based ones.
Utilizing ProbTS, we conducted a comprehensive benchmarking and analysis of a diverse range of state-of-the-art models from different strands of research. We mainly assessed these models using NAME and CRPS metrics across multiple forecasting horizons, repeating each experiment five times with different seeds to ensure result reliability.
Selected Models for Comparison.
Our selection criteria for models focused on a balance of performance, reproducibility, and simplicity. For long-term point forecasting, we included models like iTransformer [39], PatchTST [48], TimesNet [69], N-HiTS [12], and LTSF-Linear [72]. Probabilistic forecasting methods selected include GRU NVP, GRU MAF, Trans MAF [56], TimeGrad [55], and CSDI [60]. Additionally, general architectures like Linear, GRU [13], and Transformer [64], along with simple non-parametric baselines, were evaluated as a reference. For foundation models, reproducible methods such as Lag-Llama [54], TimesFM [16], Timer [40], MOIRAI [68], Chronos [3], and UniTS [24] were included. Detailed implementation specifics are in Appendix B.3.
Due to space constraints, comprehensive comparison results are placed in Appendix C, with detailed results for short-term and long-term forecasting in Tables 9 and 10, respectively. Zero-shot evaluations of pre-trained time-series foundation models are detailed in Tables 11 and 12. Our evaluation highlights the critical relationship between forecasting requirements, data properties, and modeling strategies. It aims to shed light on the strengths and limitations of current approaches, paving the way for uncovering novel research avenues.
4.1 Analyzing Classical Models for Time-series Forecasting
We examine traditional non-universal time-series models from distinct research branches: one branch focuses on developing customized neural architectures tailored for long-term point forecasting, while the other branch concentrates on creating advanced probabilistic methods for short-term distributional forecasting. Our investigation confirms the effectiveness of these models for their intended purposes. However, we observe a notable trend: the strengths of these methods tend to diminish in scenarios where they are seldom tested.
@@ 4.1 Analyzing Classical Models for Time-series Forecasting
Diminishing Advantages of Customized Architectures in Short-term Forecasting Scenarios
The comparative analysis presented in Figures 2 and 2 showcases the performance of point and probabilistic forecasting methods with respect to the NMAE metric. These figures also illustrate how NMAE values correlate with non-Gaussianity, a measure we employ to evaluate the complexity of data distributions. It becomes evident that customized architectures, originally crafted for long-term forecasting, tend to lose their competitive performance in short-term scenarios. This phenomenon could be attributed to the increased importance of accurately characterizing complex data distributions within shorter forecasting windows, where higher non-Gaussianity scores are indicative of this necessity. A closer look at Figures 2 and 2 further reveals that the performance disparity measured with CRPS becomes even more pronounced for datasets characterized by significant non-Gaussianity, such as Solar-S. This observation underscores the critical need for incorporating short-term patterns and distributional estimation capabilities into the design of new forecasting architectures.
Significant Performance Degradation for Existing Probabilistic Methods in Long-term Distributional Forecasting
The performance of current probabilistic forecasting models in long-term scenarios, even when assessed using distributional metrics such as CRPS, reveals notable limitations. This is highlighted by the comparison between Figures 2 and 2, which shows a significant drop in performance for models like TimeGrad on ETTm1-L, GRU NVP on ETTh2-L, and Weather-L datasets. The decline in performance can be attributed to the fact that these probabilistic models were not specifically designed with the unique challenges of long-term forecasting in mind. This oversight has mixed influences. On the positive side, the design of these methods has led to a more balanced approach in the choice between AR and NAR decoding schemes, providing a versatile foundation for probabilistic forecasting. However, the downside is more significant: no matter using which decoding schemes, existing probabilistic models face considerable challenges when applied to long-term distributional forecasting. We will dive deeper into the specific challenges associated with each decoding scheme next. Here the performance gap underscores the need for future research to systematically investigate long-term distributional forecasting.
Different Decoding Schemes & Challenges in Long-term Distributional Forecasting
@@
Existing probabilistic forecasting methods exhibit a balanced preference for both AR and NAR decoding schemes. For instance, TimeGrad employs an AR decoding scheme, whereas CSDI utilizes an NAR approach. This contrasts starkly with the aforementioned customized architectures, which solely opt for NAR decoding. These two types of decoding schemes, however, confront distinctive challenges when applied to long-term probabilistic forecasting. AR probabilistic models like TimeGrad tend to accumulate errors as they forecast further into the future or as the trend strength increases. This issue, which leads to a growing performance gap for longer forecasts or stronger trends, is illustrated in Figures 3 and 3. On the other hand, NAR probabilistic models, such as CSDI, face challenges related to memory usage and processing time in long-term forecasts, which are elaborated in Appendix Section D.6. Furthermore, as indicated in Table 10, even for relatively small-scale datasets, such as ETTm2 and ETTh1, CSDI’s performance in long-term scenarios is less than optimal. This suggests that its learning efficacy is also compromised by the extension of the forecasting horizon.
The Unexpected Superiority of AR Decoding in Addressing Strong Seasonality
Despite its drawbacks, AR decoding, as used in TimeGrad, excels in capturing strong seasonality, outperforming models like PatchTST in scenarios such as the Traffic dataset (Table 10). This advantage is further analyzed in Figures 3 and 3, showing AR’s increasing benefit with stronger seasonal patterns. This suggests AR’s potential in long-term forecasting could be revitalized with solutions to its error accumulation challenge in long horizons.
[NMAE w.r.t. forecasting horizon.] Refer to caption [$\frac{\text{Model.CRPS}-\text{CSDI.CRPS}}{\text{CSDI.CRPS}}$ w.r.t. non-Gaussianity.] Refer to caption
Figure 4: We evaluate the efficacy of time-series foundation models for various forecasting horizons and distributional estimation. Subplot (a), derived from Table 11 and excluding results from the Electricity dataset, demonstrates the short-term forecasting capabilities and long-term error accumulation of AR-based models. Subplot (b), draw from Table 12, investigates short-term distributional estimation, highlighting the performance challenges of foundation models compared to CSDI in handling complex data distributions. Note that we include MOIRAI with two different context lengths, 96 and 5000, as context length significantly affects its transfer performance.
4.2 Analyzing Foundation Models for Universal Time-series Forecasting
We next explore the capabilities of recent foundation models in universal time-series forecasting, focusing on their performance across different prediction horizons and their ability to estimate distributions, especially regarding their zero-shot transfer capabilities on unseen datasets. Table 11 showcases their significant progress, sometimes outperforming traditional models without re-training. Using the analytic framework of ProbTS, we delve into their methodological pros and cons.
Navigating the AR Decoding Challenge over Extended Forecasting Horizons
Figure 4 illustrates how the performance of various time-series foundation models evolves in relation to expanding forecasting horizons. It is evident that for shorter horizons, AR-based foundation models such as TimesFM and Timer exhibit highly competitive performance, on par with NAR-based models like MOIRAI. However, the advantage of NAR-based decoding becomes increasingly apparent as the forecasting horizon lengthens, as demonstrated by the widening performance gap between TimesFM and MOIRAI. This trend is consistent with our earlier observation that AR-based methods could suffer from significant error accumulation when applied to long-term time-series forecasting. Given the inherent strengths of AR decoding, such as its superiority at capturing strong seasonality and its robust performance in certain short-term forecasting scenarios, it is clear that further research is warranted to explore ways to overcome its limitations in long-term forecasting contexts. This could potentially unlock new avenues for enhancing the versatility and effectiveness of AR-based time-series foundation models across a broader range of forecasting horizons.
The Critical Role of Addressing Complex Data Distributions
Figure 4 illustrates the incremental changes in CRPS among leading probabilistic time-series foundation models, such as MOIRAI and Chronos, compared to the best-performing short-term probabilistic model, CSDI. It becomes apparent that in scenarios characterized by complex data distributions, indicated by higher non-Gaussianity scores, the performance decline of MOIRAI in relation to CSDI becomes notably more pronounced. This phenomenon may be attributed to MOIRAI’s approach to supporting distributional forecasting, which involves utilizing a mixture of predefined distribution heads. While this method is efficient and effective for certain applications, it may lack the expressiveness required to accurately model more complex data distributions. Furthermore, these observations underscore that, in specific contexts, foundation models might not be able to fully replace traditional models that have been specifically tailored and trained for particular domains. Additionally, the prospect of fine-tuning these foundation models as a remedy is less economically viable, primarily due to their significantly larger size. This highlights the importance of not only continuing to refine foundation models to enhance their adaptability and performance across a wide spectrum of data distributions but also recognizing the continued relevance of domain-specific models, especially for handling intricate data distributions where a more nuanced approach may be necessary.